Exercise 7.25

Bonita Products manufactures two component parts: AJ40 and AJ60. AJ40 components are being introduced currently, and AJ60 parts have been in production for several years. For the upcoming period, 2,600 units of each product are planned for manufacturing. Assume that the only relevant overhead cost is for engineering change orders (any requested changes in product design or the manufacturing process). AJ40 components are expected to require 4 change orders and AJ60 only 2. Each AJ40 requires 1 machine hour, and each AJ60 requires 1.5 machine hours. The cost of a change order is $560.

Estimate the cost of engineering change orders for AJ40 and AJ60 components if Bonita uses a traditional costing method and machine hours as the allocation base. (Round machine hour rate to 2 decimal places, e.g. 25.69 and final answers to 0 decimal places, e.g. 5,275.)

Total allocated cost for AJ40 $

Total allocated cost for AJ60 $

LINK TO TEXT

LINK TO TEXT

Now suppose that Bonita uses an ABC system and allocates the cost of change orders by using as cost driver the number of change orders. Estimate the cost for change orders for each unit of AJ40 and AJ60.

Total allocated cost for AJ40 $

Total allocated cost for AJ60 $

LINK TO TEXT

LINK TO TEXT

Calculate the difference in overhead allocated to each product. (Round answers to 0 decimal places, e.g. 5,275. Show an undercharge preceded by a minus sign, e.g. -256 or (256).)

Overcharge (Undercharge) for AJ40 $

Overcharge (Undercharge) for AJ60 $
Answer #1

Please find below table useful to compute desired results: -

А 1 Total Engineering change cost 2 Total allocated cost for AJ40 3 Remainder is allocated to A360 B =560*6 =(1*2600/(1*2600+

C D А 1 Total Engineering change cost 2 Total allocated cost for AJ40 3 Remainder is allocated to A360 B $3,360.00 $1,344.00